Mussels are Canada’s top shellfish aquaculture product, produced in Newfoundland and Labrador, Nova Scotia, Prince Edward Island and Quebec. Mussels filter the sea water between their shells and accumulate viral, bacterial, and parasitic pathogens in their tissues. Among the mussels collected from a certain Canadian coastline 80% are infected with a parasite. A biologist examines 120 randomly selected mussels in this region. What is the probability that more than 86% of the mussels in this sample are infected wit this parasite?
Answer #1

Solution:

Given ,

p = 80% = 0.80 (population proportion)

n = 120 (sample size)

Let \hat p be the sample proportion.

\therefore The sampling distribution of \hat p is approximately normal with

mean = \mu_{\hat p} =  p = 0.80

SD = \sigma_{\hat p}   =   p(1 - p)/n =  V0.80(1 – 0.80)/120 =  0.03651483716

Now ,

P[more than 86%]

= P[\hat p > 0.86]

=  P(-up)/o > (0.86 – Hp)/0p)

=  P(Z > (0.86-0.80)/0.03651483716)

= P(Z > 1.64)

=1 - P(Z < 1.64)

= 1 - 0.9495 ...use z table

= 0.0505

Required probability is

0.0505